The mayor of a town has proposed a plan for the construction of an adjoining community. A political study took a sample of 900 voters in the town and found that 45% of the residents favored construction. Using the data, a political strategist wants to test the claim that the percentage of residents who favor construction is over 42%. Determine the P-value of the test statistic. Round your answer to four decimal places.
Answer #1

Solution :

This is the right tailed test .

= 0.45

Test statistic = z

= - P0 / [P0 * (1 - P0 ) / n]

= 0.45 - 0.42 / [(0.42 * 0.58) / 900]

= 1.82

P(z > 1.823) = 1 - P(z < 1.823) = 0.0341

P-value = 0.0341